OpenStudy (anonymous):

what is the simplified version of sqrt(r^33)

OpenStudy (anonymous):

r^16.5

OpenStudy (anonymous):

can it be written as r^16sqrt(r) also

OpenStudy (cedric):

r^16.5

OpenStudy (turingtest):

I think they want\[\large \sqrt{r^{33}}=\sqrt{r^{32}\cdot r}=r^{16}\sqrt r\]but in general\[\large\sqrt[b]{x^a}=x^{a/b}\]so both of the above are technically correct

OpenStudy (turingtest):

so you have to be a little more specific as to what they mean by "simplify" I think it's the first one I wrote though if I had to guess